# Generating Directed Graphs with Dual Attention and Asymmetric Encoding

### v1

@@ Appendix D Directo Training and Sampling Algorithms
5:     Iteratively sample $G_{t}\sim\bm{p}_{t|1}(G_{t}|G)$triangleright Noising process
6:     $h\leftarrow\operatorname{PosEnc}(G_{t})$triangleright Positional encoding
7:     $\bm{p}_{1|t}^{\theta}(\cdot|G_{t})\leftarrow f_{\theta}(G_{t},h,t)$triangleright Denoising prediction with dual attention
@@
Alg. 2 describes the generative sampling process in which new digraphs are synthesized by iteratively denoising from an initial random digraph structure. Again, we consider $\bm{p}_{0}=\bm{p}_{noise}$ the predefined noise distribution, $f^{\theta}$ the denoising graph transformer with dual attention, $\bm{p}_{1|t}^{\theta}$ the denoising predictions, and $\bm{p}^{\theta}_{t+\Delta t|t}$ as the update rule, see Eq. 44.
@@
3:     Sample $N$ from train set $\triangleright$ # nodes
@@
6:         $h\leftarrow\operatorname{PosEnc}(G_{t})$triangleright Positional encoding
7:         $\bm{p}_{1|t}^{\theta}(\cdot|G_{t})\leftarrow f_{\theta}(G_{t},h,t)$triangleright Denoising prediction with dual attention
8:         $G_{t+\Delta t}\sim\bm{p}^{\theta}_{t+\Delta t|t}(G_{t+\Delta t}|G_{t})$triangleright Update rule
@@ E.1 Synthetic datasets
Price’s model price1965network
@@
Since the number of nodes is fixed to $N=64$, mean out-degree results in $m=6$. In practice, we implement this "bag" of nodes that replicates existing connections. For each new node $i$, we sample $m$ destination nodes from the bag. Then, for each selected node $j$, we set $\bm{A}_{ij}=1$ and add $i$ and the selected nodes to the bag. This ensures a directed acyclic graph (DAG) due to the order of node addition.
Stochastic Block Model holland1983sbm
@@ E.2 TPU Tiles
In this dataset introduced in google2023tpugraphs , each DAG represents a computation within a machine learning workload, such as a training epoch or an inference step. Each of the 12602 data points includes a computational graph, a specific compilation configuration, and the execution time of the graph when compiled with that configuration. The dataset features different model architectures, such as ResNets, EfficientNets, Masked R-CNNs, and Transformers, with graphs of up to $\sim$ 400 nodes.
@@
We adopt the processing and splits from li2024layerdag . Details on dataset statistics for the three different splits can be found in Table 4.
@@ E.3 Visual Genome
The Visual Genome dataset krishna2017visualgenome aims at bridging computer vision and natural language understanding by providing richly annotated images. It comprises over 100,000 images annotated with object labels, attributes, relationships, which allows to capture the presence of objects but also their attributes and interactions. It has become a standard benchmark for tasks such as scene graph generation, visual question answering, and grounded language understanding.
@@
Once we built the digraphs, we select a relevant subset from this raw dataset by keeping graphs with 20 to 40 nodes, and with a minimum of 25 edges per graph. Then, for each graph, we only consider objects, attributes, and relationships that are in the top-20 in terms of prevalence in the full dataset, ending with 60 node classes. We randomly split the graphs into 203 train, 52 validation and 63 test graphs, with detailed statistics of each split available in Table 5. The total number of cyclic graph in the dataset is 119 ($\sim$37.5of the digraphs).
@@ F.1 Validity metrics
Given a graph $G=(V,E)$ with $n=|V|$ nodes and $m=|E|$ edges, we want to determine whether the graph likely originates from an Erdős–Rényi model $G(n,p)$ with edge probability $p$ using a Wald test held2013appliedstats . For that, we follow the following steps distinguishing the cases in which the graphs are DAGs or not:
@@
    $m_{\text{max}}=\begin{cases}\frac{n(n-1)}{2}&\text{if the graph acyclic}\\ n(n-1)&\text{if the graph is not acyclic}\end{cases}$
@@
    Block assignment via model inference: Given the adjacency matrix $A\in\{0,1\}^{n\times n}$ of a graph $G$, we use an algorithm peixoto2014sbminference to infer the stochastic block model (block assignment) from a given network:
    Infer $\text{Infer }z:V\to\{1,\dots,B\}\quad\text{using }\min\mathcal{L}(G,z)$
@@ F.3 Maximum Mean Discrepancy metrics
For the different metrics, we adapt previous work from martinkus2022SPECTRE ; law2024directedheatkernels . We propose to measure both out-degree and in-degree distributions, as well as clustering, spectre and walvelet. To measure the clustering coefficient, we compute the distribution of directed local clustering coefficients using networkx function clustering(), which supports digraphs. For the spectral features (spectre and wavelet), we derive the spectral features from the directed Laplacian, as described in Section 4.
In addition, the orbit metric that was also computed in martinkus2022SPECTRE has not been adapted to the directed setting as it relies on the Orbit Counting Algorithm (ORCA) which counts graphlets in networks but is only implemented in the undirected setting. Therefore, adapting this metric to the directed setting remains an open challenge.
@@ G.1 Details on baselines
    $P(n)=\frac{\text{Number of training graphs with }n\text{ nodes}}{\text{Total number of training graphs}}.$
@@
    $P(v=c)=\frac{\text{Number of nodes of class }c}{\text{Total number of nodes inall training graphs}}.$
@@
    $P(e_{ij}=k\mid v_{i}=c_{a},v_{j}=c_{b})=\frac{\text{Number of edges of type }k\text{ between }(c_{a},c_{b})}{\sum_{k^{\prime}=1}^{K}\text{Number of edges oftype }k^{\prime}\text{ between }(c_{a},c_{b})}.$
@@ G.2 Training setup
For RRWP, we set the walk length parameter to $K=20$. For MagLap with $Q=5$, we used equidistant potentials $\bm{q}_{5}=(0,0.1,0.2,0.3,0.4)$. For $Q=10$, we selected again 10 equidistant potentials $\bm{q}_{10}=(0.01,\ldots,0.1)$. In all cases, we then kept the $k=10$ first eigenvalues and eigenvectors (padding with null values whenever the graphs had less than 10 nodes).
@@
For Directo-DD we performed 500 sampling steps in all the different model configurations.
@@
Baseline Revcos 0.3 5
RRWP Polyinc 0.01 10
@@
Baseline Cos 0.05 0
RRWP Polydec 0.3 100
MagLap Polydec 0.1 10
@@ G.3 Resources and runtime
Baseline 120 13.6 40 0.125
RRWP 120 13.5 40 0.3
MagLap 120 14.4 40 0.8
@@
Baseline 120 23.7 40 0.3
RRWP 120 13.7 40 0.5
MagLap 120 21 40 2.1
@@ Appendix H Additional results
In this section, we present additional results that offer deeper insights into our model’s performance and design choices. Specifically, we analyze the impact of the dual attention mechanism and the choice of positional encoding, compare performance on ER and DAG accuracy metrics, examine how dataset size influences generation quality, and evaluate the effect of the sampling optimization step in discrete flow matching.
H.1 The role of dual attention

### v2

@@ Appendix D Directo Training and Sampling Algorithms
5:  Iteratively sample $G_{t}\sim\bm{p}_{t|1}(G_{t}|G)$ $\triangleright$ Noising process
6:  $h\leftarrow\operatorname{PosEnc}(G_{t})$ $\triangleright$ Positional encoding
7:  $\bm{p}_{1|t}^{\theta}(\cdot|G_{t})\leftarrow f_{\theta}(G_{t},h,t)$ $\triangleright$ Denoising prediction with dual attention
@@
Alg. 2 describes the generative sampling process in which new digraphs are synthesized by iteratively denoising from an initial random digraph structure. Again, we consider $\bm{p}_{0}=\bm{p}_{noise}$ the predefined noise distribution, $f^{\theta}$ the denoising graph transformer with dual attention, $\bm{p}_{1|t}^{\theta}$ the denoising predictions, and $\bm{p}^{\theta}_{t+\Delta t|t}$ as the update rule, see Equation 41.
@@
3:  Sample NN from train set $\triangleright$ # nodes
@@
6:   $h\leftarrow\operatorname{PosEnc}(G_{t})$ $\triangleright$ Positional encoding
7:   $\bm{p}_{1|t}^{\theta}(\cdot|G_{t})\leftarrow f_{\theta}(G_{t},h,t)$ $\triangleright$ Denoising prediction with dual attention
8:   $G_{t+\Delta t}\sim\bm{p}^{\theta}_{t+\Delta t|t}(G_{t+\Delta t}|G_{t})$ $\triangleright$ Update rule
@@ E.1 Synthetic datasets
Price’s model (de Solla Price, 1965)
@@
Since the number of nodes is fixed to $N=64$, mean out-degree results in $m=6$. In practice, we implement this ”bag” of nodes that replicates existing connections. For each new node $i$, we sample $m$ destination nodes from the bag. Then, for each selected node $j$, we set $\bm{A}_{ij}=1$ and add $i$ and the selected nodes to the bag. This ensures a directed acyclic graph (DAG) due to the order of node addition.
Stochastic Block Model (Holland et al., 1983)
@@ E.2 TPU Tiles
In this dataset introduced by Phothilimthana et al. (2023), each DAG represents a computation within a machine learning workload, such as a training epoch or an inference step. Each of the 6301 datapoints includes a computational graph, a specific compilation configuration, and the execution time of the graph when compiled with that configuration. The dataset features different model architectures, such as ResNets, EfficientNets, Masked R-CNNs, and Transformers, with graphs of up to $\sim$ 400 nodes.
@@
We adopt the processing and splits from Li et al. (2025). Details on dataset statistics for the three different splits can be found in Table 4.
@@ E.3 Visual Genome
The Visual Genome dataset (Krishna et al., 2017) aims at bridging computer vision and natural language understanding by providing richly annotated images. It comprises over 100,000 images annotated with object labels, attributes, relationships, which allows to capture the presence of objects but also their attributes and interactions. It has become a standard benchmark for tasks such as scene graph generation, visual question answering, and grounded language understanding.
@@
Once we built the digraphs, we select a relevant subset from this raw dataset by keeping graphs with 20 to 40 nodes, and with a minimum of 25 edges per graph. Then, for each graph, we only consider objects, attributes, and relationships that are in the top-20 in terms of prevalence in the full dataset, ending with 60 node classes. We randomly split the graphs into 203 train, 52 validation and 63 test graphs, with detailed statistics of each split available in Table 5. The total number of cyclic graph in the dataset is 119 ($\sim$37.5% of the digraphs).
@@ F.1 Validity metrics
Given a graph $G=(V,E)$ with $n=|V|$ nodes and $m=|E|$ edges, we want to determine whether the graph likely originates from an Erdős–Rényi model $G(n,p)$ with edge probability $p$ using a Wald test (Held & Sabanés Bové, 2013). For that, we follow the following steps distinguishing the cases in which the graphs are DAGs or not:
@@
[…]
@@
    Block assignment via model inference: Given the adjacency matrix $A\in\{0,1\}^{n\times n}$ of a graph $G$, we use an algorithm (Peixoto, 2014) to infer the stochastic block model (block assignment) from a given network:
    $\text{Infer }z:V\to\{1,\dots,B\}\quad\text{using }\min\mathcal{L}(G,z)$
@@ F.3 Maximum Mean Discrepancy metrics
For the different metrics, we adapt previous work from Martinkus et al. (2022); Law et al. (2025). We propose to measure both out-degree and in-degree distributions, as well as clustering, spectre and walvelet. To measure the clustering coefficient, we compute the distribution of directed local clustering coefficients using networkx function clustering(), which supports digraphs. For the spectral features (spectre and wavelet), we derive the spectral features from the directed Laplacian, as described in Section 4.
In addition, the orbit metric that was also computed in Martinkus et al. (2022) has not been adapted to the directed setting as it relies on the Orbit Counting Algorithm (ORCA) which counts graphlets in networks but is only implemented in the undirected setting. Therefore, adapting this metric to the directed setting remains an open challenge.
F.4 Joint node-edge distributional metrics
Downstream evaluations in graph generative models usually assume conditional graph generation and alignment with auxiliary inputs. Moreover, standard evaluations in domains such as Scene Graph Generation (Li et al., 2024; Lorenz et al., 2024) assume access to input images to compare generated graphs against ground truth images using retrieval metrics such as Precision@K and Recall@K. These require alignment with specific input conditions, which is not possible in our formulation as there are no ground truth images for the generated graphs. Instead, we propose metrics to directly evaluate generative performance by assessing joint node-edge distributional coverage, an essential aspect for labeled graphs such as NAS and scene graphs.
  • Node type distribution MMD: To evaluate coverage of node class distributions, we compute normalized histograms over node labels for each graph and calculate the MMD score between generated and reference sets, which captures how well categorical proportions are preserved across generated graphs.
  • Triplet-based precision and recall: Inspired by retrieval-style evaluations in Scene Graph Generation (Li et al., 2024; Lorenz et al., 2024), we extract semantic tuples from graphs. For scene graphs, we use $(\text{subject},\text{attribute})$ pairs and $(\text{object},\text{relation},\text{object})$ triplets. For the TPU Tiles dataset, we extract element–element pairs. We compare the tuples from generated graphs with those in the test set and compute:
    $\text{Precision}=\frac{|\text{correctly generated tuples}|}{|\text{generated tuples}|}\quad\text{Recall}=\frac{|\text{correctly generated tuples}|}{|\text{test tuples}|}$
  • Embedding-based distances (FID and RBF): Following Thompson et al. (2022), we compute RBF-kernel MMD and Fréchet Inception Distance (FID) over embeddings extracted from a random GNN trunk. These embeddings encode both node labels and structural information, thus capturing joint node-edge distributions.
F.5 Downstream tasks metrics
In addition to the distributional alignment evaluations, we also provide a evaluations on downstream tasks for the TPU Tiles dataset. Here, we train a conditional generative model and follow the evaluation protocol of LayerDAG (Li et al., 2025). Specifically, we employ their ML-based surrogate cost models to approximate the computational performance of generated architectures. This ensures direct comparability with prior work.
@@ G.1 Details on baselines
[…]
@@
[…]
@@
[…]
@@ G.2 Training setup
For RRWP, we set the walk length parameter to $K=20$. For MagLap with $Q=5$, we used equidistant potentials $\bm{q}_{5}=(0,0.1,0.2,0.3,0.4)$. For $Q=10$, we selected again 10 equidistant potentials $\bm{q}_{10}=(0.01,\ldots,0.1)$. In all cases, we then kept the $k=10$ first eigenvalues and eigenvectors, padding with null values whenever the graphs had less than 10 nodes.
Conditional generation
To train with classifier-free guidance conditioning, we first conducted a hyperparameter search to determine the guidance strength with values from 0 to 4, in particular $\omega_{cond}\in(0,0.25,0.5,1,1.5,2,3,4)$. We ultimately set the parameter to $\omega_{cond}=0.25$.
@@
Finally, for the ablations on dual attention and positional encodings, we did not perform sampling optimization on $\eta$ and $\omega$ to avoid the consequent computational cost, but searched for the optimal time distortion. For the scalability experiments, we also did not perform sampling optimization but chose the relevant optimal configuration (from Table 7). For Directo-DD we performed 500 sampling steps in all the different model configurations.
@@
Baseline Revcos 0 0
RRWP Polyinc 0.1 50
@@
Baseline Cos 0.1 0
RRWP Polydec 0.3 10
MagLap Polydec 0.1 50
@@ G.3 Resources and runtime
Baseline 128 13.6 40 0.125
RRWP 128 13.5 40 0.3
MagLap 128 14.4 40 0.8
@@
Baseline 128 23.7 40 0.3
RRWP 128 13.7 40 0.5
MagLap 128 21 40 2.1
@@ Appendix H Additional results
In this section, we present additional results that offer deeper insights into our model’s performance and design choices. Specifically, we analyze the impact of the dual attention mechanism and the choice of positional encoding, compare performance on ER and DAG accuracy metrics, examine how dataset size influences generation quality, provide detailed results for conditional generation, and evaluate the effect of the sampling optimization step in discrete flow matching.
H.1 Extended results for synthetic datasets
Table 9 reports the performance of Directo across different configurations on synthetic datasets. We observe that the choice of architecture and positional encodings leads to noticeable differences in the distributional alignment metrics (MMD), highlighting the strong performance of Directo. At the same time, uniqueness and novelty remain consistently at 100% across all models, confirming that the V.U.N. ratio is driven uniquely by the validity metric.
Erdős-Renyi Directed Acyclic Graph (ER-DAG)
Training set 0.0113 0.0103 0.0355 0.0038 0.0024 1.0 99.2 100 0.0 0.0
MLE 0.0083 $\pm$ 0.0004 0.0089 $\pm$ 0.0003 0.1318 $\pm$ 0.0077 0.0823 $\pm$ 0.0013 0.1162 $\pm$ 0.0018 15.1 $\pm$ 0.2 0.0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
D-VAE 0.6158 $\pm$ 0.0160 0.6246 $\pm$ 0.0103 1.0509 $\pm$ 0.00304 0.6160 $\pm$ 0.0315 0.5432 $\pm$ 0.0389 106.6 $\pm$ 5.4 0.0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
DiGress 0.0138 $\pm$ 0.0035 0.0143 $\pm$ 0.0050 0.1074 $\pm$ 0.0090 0.0073 $\pm$ 0.0017 0.0042 $\pm$ 0.0011 1.9 $\pm$ 0.3 34.0 $\pm$ 4.1 100 $\pm$ 0.0 100 $\pm$ 0.0 34.0 $\pm$ 4.1
DeFoG 0.0407 $\pm$ 0.0008 0.0041 $\pm$ 0.0011 0.0460 $\pm$ 0.0052 0.0061 $\pm$ 0.0008 0.0008 $\pm$ 0.0003 1.6 $\pm$ 0.2 75 $\pm$ 2.2 100 $\pm$ 0.0 100 $\pm$ 0.0 75.0 $\pm$ 2.2
Directo-DD RRWP 0.0142 $\pm$ 0.0040 0.0129 $\pm$ 0.0034 0.0408 $\pm$ 0.0053 0.0055 $\pm$ 0.0010 0.0048 $\pm$ 0.0014 1.4 $\pm$ 0.3 79.0 $\pm$ 3.7 100 $\pm$ 0.0 100 $\pm$ 0.0 79.0 $\pm$ 3.7
Directo-DD MagLap 0.0145 $\pm$ 0.0040 0.0134 $\pm$ 0.0033 0.0582 $\pm$ 0.0083 0.0063 $\pm$ 0.0015 0.0034 $\pm$ 0.0011 1.5 $\pm$ 0.2 85.0 $\pm$ 9.2 100 $\pm$ 0.0 100 $\pm$ 0.0 85.0 $\pm$ 9.2
Directo RRWP 0.0107 $\pm$ 0.0019 0.0104 $\pm$ 0.0012 0.1209 $\pm$ 0.0194 0.0054 $\pm$ 0.0005 0.0043 $\pm$ 0.0008 1.7 $\pm$ 0.1 94.0 $\pm$ 1.0 100 $\pm$ 0.0 100 $\pm$ 0.0 94.0 $\pm$ 1.0
Directo MagLap 0.0117 $\pm$ 0.0014 0.0110 $\pm$ 0.0015 0.0711 $\pm$ 0.0120 0.0055 $\pm$ 0.0016 0.0026 $\pm$ 0.0004 1.3 $\pm$ 0.2 92.0 $\pm$ 3.7 100 $\pm$ 0.0 100 $\pm$ 0.0 92.0 $\pm$ 3.7
Stochastic Block Model (SBM)
Training set 0.0031 0.0031 0.0274 0.0027 0.0011 1.0 97.7 100 0.0 0.0
MLE 0.0020 $\pm$ 0.0002 0.0020 $\pm$ 0.0002 0.1973 $\pm$ 0.0162 0.0087 $\pm$ 0.0003 0.0508 $\pm$ 0.0009 11.6 $\pm$ 0.2 0.0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
DeFoG 0.0026 $\pm$ 0.0019 0.0022 $\pm$ 0.0016 0.0813 $\pm$ 0.0058 0.0048 $\pm$ 0.0002 0.0110 $\pm$ 0.0019 4.3 $\pm$ 0.8 37 $\pm$ 6.6 100 $\pm$ 0.0 100 $\pm$ 0.0 37.0 $\pm$ 6.6
Directo RRWP 0.0031 $\pm$ 0.0014 0.0028 $\pm$ 0.0013 0.0594 $\pm$ 0.0027 0.0035 $\pm$ 0.0004 0.0027 $\pm$ 0.0009 1.8 $\pm$ 0.5 99.5 $\pm$ 1.0 100 $\pm$ 0.0 100 $\pm$ 0.0 99.5 $\pm$ 1.0
H.2 Extended results for real-world datasets
Table 10 reports the performance of Directo across different configurations on real-world datasets. While some configurations show evidence of partial memorization, particularly in node and edge distributions, this does not negatively affect the overall generative quality: V.U.N. remains high, and Directo consistently achieves the best or second-best performance across the majority of metrics. These results highlight that, despite occasional replication of training structures, the model maintains strong distributional alignment and generates diverse, valid, and novel graphs.
TPU Tiles
Training set 0.0003 0.0003 0.0007 0.0006 0.0002 1.0 100 100 0.0 0.0
D-VAE OOT OOT OOT OOT OOT OOT OOT OOT OOT OOT
DeFoG 0.0099 $\pm$ 0.0012 0.0794 $\pm$ 0.0023 0.0042 $\pm$ 0.0027 0.0040 $\pm$ 0.0005 0.0017 $\pm$ 0.0003 63.7 $\pm$ 2.6 86.3 $\pm$ 2.2 87.7(∗) $\pm$ 5.4 93.3 $\pm$ 2.5 72.0 $\pm$ 2.4
Directo-DD MagLap 0.0115 $\pm$ 0.0035 0.0703 $\pm$ 0.0033 0.0103 $\pm$ 0.0021 0.0076 $\pm$ 0.0007 0.0043 $\pm$ 0.0014 64.3 $\pm$ 5.3 86.5 $\pm$ 5.1 90.5 $\pm$ 3.3 100 $\pm$ 0.0 77.0 $\pm$ 7.0
Directo RRWP 0.0133 $\pm$ 0.0025 0.0859 $\pm$ 0.0072 0.0136 $\pm$ 0.0075 0.0086 $\pm$ 0.0010 0.0038 $\pm$ 0.0009 75.4 $\pm$ 8.1 97.0 $\pm$ 1.0 81.8(∗) $\pm$ 4.5 96.5 $\pm$ 1.2 77.0 $\pm$ 2.9
Visual Genome
Training set 0.0018 0.0030 0.0000 0.0072 0.0036 1.0 100 100 0.0 0.0
MLE 0.0607 $\pm$ 0.0036 0.0474 $\pm$ 0.0023 0.5342 $\pm$ 0.0830 0.0535 $\pm$ 0.0015 0.0399 $\pm$ 0.0017 17.0 $\pm$ 0.6 0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0.0 $\pm$ 0.0
Directo-DD RRWP 0.0424 $\pm$ 0.0043 0.0413 $\pm$ 0.0037 0.0000 $\pm$ 0.0000 0.0132 $\pm$ 0.0053 0.0074 $\pm$ 0.0008 15.3 $\pm$ 0.8 72.7 $\pm$ 3.9 100 $\pm$ 0.0 100 $\pm$ 0.0 72.7 $\pm$ 3.9
Directo-DD MagLap 0.0245 $\pm$ 0.0022 0.0359 $\pm$ 0.0036 0.0000 $\pm$ 0.0000 0.0163 $\pm$ 0.0026 0.0086 $\pm$ 0.0011 7.6 $\pm$ 0.7 86.5 $\pm$ 5.1 100 $\pm$ 0.0 100 $\pm$ 0.0 61.9 $\pm$ 4.4
Directo RRWP 0.0494 $\pm$ 0.0021 0.0425 $\pm$ 0.0029 0.0000 $\pm$ 0.0000 0.0226 $\pm$ 0.0075 0.0228 $\pm$ 0.0058 12.8 $\pm$ 0.6 86.0 $\pm$ 4.5 98.4 $\pm$ 1.7 99.3 $\pm$ 0.7 83.8 $\pm$ 4.3
Directo MagLap 0.0180 $\pm$ 0.0024 0.0302 $\pm$ 0.0040 0.0000 $\pm$ 0.0000 0.0142 $\pm$ 0.0021 0.0099 $\pm$ 0.0013 6.2 $\pm$ 0.5 67.6 $\pm$ 3.6 99.7 $\pm$ 0.6 99.7 $\pm$ 0.6 67.0 $\pm$ 4.3
Table 11 reports the node-label distributional alignment metrics for the two real-world datasets considered. The results highlight that Directo consistently manages to capture node-label distributions accurately. On TPU Tiles, Directo achieve the lowest node-type MMD, indicating strong alignment with the training set, while also maintaining high precision and competitive recall. On Visual Genome, although Directo exhibits slightly lower node-type MMD compared to DeFoG, it attains the highest precision, recall, and RBF MMD, reflecting its ability to capture the underlying distribution. Overall, these results indicate that Directo effectively balances reproducing node label distributions while capturing richer graph structure, with the different positional encoding strategies providing complementary strengths in distributional alignment.
It is relevant to note that the low precision achieved in the training sets is due to the fact that the ground truth set is much larger than the generated set, and therefore there are many pairs and triplets present in the ground truth dataset that have not been generated.
Table 11: Directed graph generation performance on node-label distributional alignment metrics for the two real-world datasets. Results are presented as mean $\pm$ standard deviation across five sampling runs.
Model Node type (MMD) $\downarrow$ Precision $\uparrow$ Recall $\uparrow$ FID $\downarrow$ RBF (MMD) $\downarrow$
TPU Tiles
Training set 0.0001 0.64 0.99 0.0854 0.0021
MLE 0.4573 $\pm$ 0.0070 0.10 $\pm$ 0.01 0.22 $\pm$ 0.01 105776.8 $\pm$ 22676.3 1.0392 $\pm$ 0.0326
D-VAE OOT OOT OOT OOT OOT
LayerDAG 0.0130 $\pm$ 0.0077 0.31 $\pm$ 0.03 0.24 $\pm$ 0.10 2864.1 $\pm$ 1715.6 1.0208 $\pm$ 0.0225
DiGress 0.0037 $\pm$ 0.0033 0.95 $\pm$ 0.03 0.46 $\pm$ 0.05 887.7 $\pm$ 313.8 0.0969 $\pm$ 0.0331
DeFoG 0.0031 $\pm$ 0.0015 0.90 $\pm$ 0.03 0.45 $\pm$ 0.05 159.5 $\pm$ 108.9 0.0579 $\pm$ 0.0150
Directo-DD RRWP 0.0028 $\pm$ 0.0008 0.92 $\pm$ 0.03 0.48 $\pm$ 0.02 336.6 $\pm$ 163.9 0.0587 $\pm$ 0.0225
Directo-DD MagLap 0.0051 $\pm$ 0.0030 0.91 $\pm$ 0.03 0.42 $\pm$ 0.03 482.2 $\pm$ 124.3 0.0792 $\pm$ 0.0266
Directo RRWP 0.0018 $\pm$ 0.0004 0.97 $\pm$ 0.01 0.44 $\pm$ 0.03 293.0 $\pm$ 110.1 0.0438 $\pm$ 0.0179
Directo MagLap 0.0086 $\pm$ 0.0014 0.96 $\pm$ 0.02 0.42 $\pm$ 0.02 96.0 $\pm$ 64.0 0.0421 $\pm$ 0.0080
Visual Genome
Training set 0.0072 0.34 0.81 1.2 0.0214
MLE 0.0807 $\pm$ 0.0030 0.25 $\pm$ 0.03 0.13 $\pm$ 0.01 48941.0 $\pm$ 19586.5 0.6179 $\pm$ 0.0253
DiGress 0.0206 $\pm$ 0.0034 0.46 $\pm$ 0.04 0.35 $\pm$ 0.04 42.6 $\pm$ 1.3 0.2321 $\pm$ 0.0276
DeFoG 0.0181 $\pm$ 0.0030 0.51 $\pm$ 0.02 0.39 $\pm$ 0.02 19.9 $\pm$ 5.6 0.0845 $\pm$ 0.0233
Directo-DD RRWP 0.0231 $\pm$ 0.0082 0.48 $\pm$ 0.02 0.48 $\pm$ 0.02 4.8 $\pm$ 1.2 0.0391 $\pm$ 0.0047
Directo-DD MagLap 0.0288 $\pm$ 0.0069 0.50 $\pm$ 0.02 0.46 $\pm$ 0.02 8.4 $\pm$ 2.5 0.0420 $\pm$ 0.0056
Directo RRWP 0.0272 $\pm$ 0.0056 0.53 $\pm$ 0.02 0.45 $\pm$ 0.02 4.7 $\pm$ 2.1 0.0383 $\pm$ 0.0049
Directo MagLap 0.0299 $\pm$ 0.0047 0.58 $\pm$ 0.02 0.46 $\pm$ 0.01 9.3 $\pm$ 2.6 0.0507 $\pm$ 0.0118
H.3 The role of dual attention